Grade 12th passPhysical Chemistry

one mole of water at 100 degree celcius is converted into steam at 100 degree celcius at constant pressure of 1 atm. The change in entropy is (heat of vaporisation of water at 100 degree celcius is 540 cal/gm)

Arisha Ali Rahi
8 Years agoGrade 12th pass

1 Answer

Eshan
8 Years ago
Dear student,

Total mass of 1 mole of water=18gm

Heat required to vaporise this amount of water=mL_V=540\times 18cal=40.63kJ

Hence change in entropy=\dfrac{\Delta Q}{T}=\dfrac{40.63\times 10^3}{373K}\approx 109J/K